OpenStudy (anonymous):

Calculus: Find the derivative. y=x^3cosx

OpenStudy (anonymous):

\[x ^{3}\cos x\]

OpenStudy (anonymous):

product rule \[(f(x)g(x))^{\prime}=f^{\prime}(x)g(x)+f(x)g^{\prime}(x)\]

OpenStudy (anonymous):

can u explain more pls? xD

OpenStudy (anonymous):

like how to do it cause i dont get it :\

OpenStudy (anonymous):

\[f(x)=x^3;g(x)=\cos{x};\;\; [f(x)g(x)]^{\prime} = (x^3)^{\prime}\cos{x}+x^3(\cos{x})^{\prime}\]

OpenStudy (anonymous):

Do you not understand how to do the actual derivatives or the product rule?

OpenStudy (anonymous):

I just dont know how to do this problem :)

OpenStudy (anonymous):

\[(x^3)^{\prime} = 3x^{(3-1)}\] from the power rule \[(\cos{x})^{\prime} = -\sin{x}\] \[3x^2\cos{x}-x^3\sin{x}\]

OpenStudy (anonymous):